# Omni-Router: Sharing Routing Decisions in Sparse Mixture-of-Experts for Speech Recognition

### v2

@@ Abstract
Index Terms:
Speech Recognition, Mixture-of-Experts (MoE), Omni-router Transformer
I Introduction
Automatic speech recognition (ASR) systems have made remarkable progress in recent years, driven by advances in neural network architectures and the availability of large-scale datasets [1, 2, 3, 4]. Despite these advancements, achieving high recognition accuracy across diverse acoustic and linguistic conditions remains a challenging problem, particularly given the variability in speaker accents, background noise, and domain-specific vocabularies.
There are several ways to improve ASR performance. Combining an acoustic model with additional language models (LMs), such as $n$-gram LMs [5, 6], neural LMs [5, 6, 7, 8] and error correction LMs [9, 10], were found to be effective ways to produce better transcriptions. Other approaches are inspired by the development of the scaling laws [11, 12, 13], where scaling up model and data size is crucial to improve model performance. However, classical large models often face trade-offs between computational efficiency and representational power, especially when deployed in real-world scenarios where resource constraints and robustness are critical considerations [14, 15].
Mixture-of-experts (MoE) models [16] have emerged as a promising architecture for their ability to scale model capacity (in terms of number of parameters) without incurring proportional overhead at inference. Unlike traditional dense models, which activate all parameters for every input, MoE models dynamically route input data to a subset of specialized experts tailored to process particular characteristics of the data. This routing mechanism enables MoE models to scale to significantly larger parameter counts without proportional increases in inference cost, offering both efficiency and flexibility. ASR involves highly diverse speech signals, with differences in pronunciation, accent, and noise patterns and MoE’s, by design, can be quite appropriate to modeling such diversity. By assigning specific experts to different aspects of the input, MoE models could better capture this variability, potentially improving recognition performance and robustness. Moreover, their ability to adaptively allocate resources makes them well-suited for deployment in heterogeneous environments, from edge devices with constrained resources to cloud-based systems with ample inference capacity. Recently, [17] showed that increasing the number of experts (while fixing the number of active parameters) increases the memorization performance consistently while the reasoning capabilities saturate; [18] showed interaction between compute per example and total number of parameters, and their impact on model in MoEs: there is an optimal level of sparsity that improves both training efficiency and model performance.
@@ I Introduction
The effectiveness of MoE architectures heavily relies on the routing mechanism that directs inputs to appropriate experts. Conventional MoE methods learn both the routers and expert parameters at the same time, with stochastic gradient descent (SGD). The learning is complicated by issues such as uneven expert utilization [19], insufficient expert specialization [15], and high routing computational costs [20] often resulting in training instability [21]. While these issues are well explored, in this paper we highlight an under-explored issue - namely that the correlation between which experts are used at different layers is weak, with different layers making seemly arbitrary, independent decisions about which experts to use. We hypothesize that this would lead to models that do not specialize to data very strongly. To address this issue we propose to use a shared router across all layers that encourages coordinated decision-making across MoE layers (see architecture in Fig. 1). We think it is possible to have a shared router across different layers, because of the residual nature of transformers - features from each layer are modified by residual connections and only change minimally from layer to layer [22], and thus sharing a decision boundaries across layers is possible by just using a shared router. Empirical results across multiple model configurations consistently indicate that Omni-router delivers superior lower training error, test word error rates (WER) and robustness against highly diverse data compared to traditional dense models and Switch Transformer baselines across diverse out-of-domain evaluation datasets.
II Related work
@@
MoE models were first introduced by Jacobs et al.[23] as a dynamic mechanism to combine multiple specialized sub-models through gating functions, with subsequent hierarchical extensions presented by Jordan and Jacobs[24]. Modern advances have significantly scaled MoE architectures to efficiently accommodate extremely large model sizes, exemplified by sparsely gated MoE layers proposed by Shazeer et al [16]. Further efficiency improvements were realized with the Switch Transformer[25], simplifying the routing mechanism by activating only one expert per input token. Additional strategies, such as BASE layers [26] and hash-based routing techniques [27], have been explored to mitigate issues related to load imbalance and training inefficiency. MoE architectures have demonstrated substantial success across various domains, including natural language modeling (e.g., GLAM [28]) and computer vision tasks [29].
@@
In contrast, MoE applications within ASR remain comparatively limited. SpeechMoE [30] proposes a shared embedding network and concatenates the shared embedding with output of the previous layer as the input of routers, meanwhile adding more loss terms, such as $L_{1}$, mean importance and embedding losses. Building on this work, SpeechMoE2 [31] proposes a new router architecture by integrating additional global domain and accent embedding into router input to promote adaptability. Both SpeechMoE and SpeechMoE2 predominantly focus on monolingual speech recognition tasks.
@@
Recently, the application of MoE architectures to multilingual ASR has garnered increasing attention [32, 33, 34]. Hu et al [32] integrate MoE layers within conformer architectures [3], demonstrating an average relative improvement in WER of 11.9% across twelve languages. Kwon et al.[33] propose a MoE model consisting of language-specific experts alongside a language-agnostic expert, while Li et al.[34] utilize routers that select optimal expert subsets combined with joint Connectionist Temporal Classification (CTC)/Attention training [35] and rescoring, achieving substantial improvements in low-resource multilingual ASR scenarios.
@@
Despite these advancements, critical challenges persist, particularly concerning load balancing, model stability, and training robustness [21]. Such issues motivate the exploration of novel routing mechanisms aimed at enhancing MoE model robustness and effectiveness. Our contributions address these gaps in two primary ways. First, we introduce an innovative shared-routing approach within Transformer-based MoE ASR models, demonstrating its effectiveness in terms of both accuracy and robustness compared to conventional dense and Switch Transformer MoE architectures. Second, our proposed MoE design remains straightforward, eliminating the need for auxiliary embedding networks, complex auxiliary loss combinations [30, 31], or explicit inductive biases such as accent or language embeddings [31, 33], thus facilitating simpler implementation and training while being effective.
@@
III Background
III-A Mixture-of-Experts layer
Most works consider MoE in replacement of the feed-forward network (FFN) layer in transformer models [1]. Typically, an MoE layer contains a routing network and a branch of experts, which are also FFNs. The router dispatches the incoming $T$ tokens $X^{l}\in\mathbb{R}^{T\times D}$ at layer $l$ to different experts first by estimating the weights $P\in\mathbb{R}^{T\times N}$ for each of $N$ experts:
$P^{l}(X^{l})=\mathrm{Softmax}(X^{l}W^{l}),\qquad l=1,...,L,$ (1)
@@ III-A Mixture-of-Experts layer
III-B Switch Transformer
Instead of choosing top-$k$ experts based on the router’s decision, Switch Transformer [25] simplifies the routing mechanism and chooses only the top-1 expert. It was shown that switch layer not only reduces the router computation cost but also gives better performance in language models. Throughout this paper, we also use only the top-1 expert.
III-C Load balancing loss
During training of MoE models, an auxiliary loss is often introduced to maintain the balance of load (the number of tokens that go through one expert) of each expert. As introduced in [25], given $N$ experts and $T$ tokens $X^{l}$ for layer $l$, the fraction of tokens dispatched to each expert $j$, $f^{l}$, and the fraction of router probability allocated for each expert $j$, $\rho^{l}$, are defined as
$f_{j}^{l}=\frac{1}{T}\sum_{i=1}^{T}\mathbbm{1}\{\arg\max_{j^{\prime}}[P^{l}(X^{l})]_{i}^{j^{\prime}}=j\},$ (3)
$\rho^{l}_{j}=\frac{1}{T}\sum_{i:\arg\max_{j^{\prime}}[P^{l}(X^{l})]_{i}^{j^{\prime}}=j}[P^{l}(X^{l})]^{j}_{i},$ (4)
@@ III-C Load balancing loss
$\mathcal{L}^{l}_{\mathrm{load}}=N\sum_{j=1}^{N}f_{j}^{l}\cdot\rho_{j}^{l}.$ (5)
IV Omni-router: sharing routing decisions across MoE layers
@@
$P^{l}(X^{l})=\mathrm{Softmax}(X^{l}W^{\mathrm{shared}}).$ (6)
This strategy is viable due to the prevalent use of pre-layer normalization in transformer architectures, which incorporates residual connections directly linking input and output representations. Consequently, feature representations between consecutive layers remain sufficiently similar, enabling shared routing parameters to establish coherent—though not identical—decision boundaries across layers. Meanwhile, the dense parameters within each MoE layer remain free to specialize independently, adapting to layer-specific decision boundaries. This approach contrasts with existing parameter-sharing methods in MoE architectures, which typically involve sharing the experts’ weights rather than the routing parameters themselves [36, 37, 38].
V Empirical setup
V-A Data
We utilize a large-scale conversational audio dataset collected from publicly accessible sources, named SpeechCrawl. This dataset was specifically curated to be diverse, conversational, multilingual, and multi-speaker. Audio samples average approximately 30 minutes in duration, with around 60% consisting of English speech. Since SpeechCrawl does not provide transcriptions, we generate pseudo-labels using the WhisperX [39] pipeline in combination with Whisper’s large-v2 multilingual model [2]. These pseudo-labels simultaneously yield audio segmentation (with segments shorter than 30 seconds) and corresponding transcripts. To ensure quality, we filter out unreliable pseudo-labels resulting from looping artifacts of decoding in the sequence-to-sequence model. Specifically, we employ a simple heuristic based on expected speech rates (approximately three words per second), similar to [40]. After filtering, we retain approximately 1M hours of English speech segments for training. Additionally, we randomly select about 12h of audio (24 files) as our development set to monitor training progress and identify the optimal checkpoint for subsequent evaluation on out-of-domain (OOD) datasets. For OOD testing, we employ diverse test sets from AMI-IHM [41], Callhome [42], Chime6 [43], CommonVoice [44], Fleurs [45], Librispeech [46], Switchboard [42], Tedlium [47], Voxpopuli [48], and Wall Street Journal [49]. Importantly, none of the selected test datasets overlap with the training data.
V-B ASR models
We compare our proposed Omni-router Transformer model against two baseline architectures: a dense transformer and a Switch Transformer. All ASR models utilize the standard Transformer encoder architecture [1], optimized using CTC [50] and employing an 8k-token word-piece vocabulary derived from the training transcriptions.
@@ V-B ASR models
V-C Training setup
All ASR models take as input 80-dimensional log-mel filterbank features computed using a 25ms sliding window with a 10ms stride. Subsequently, every four consecutive frames are stacked before being fed into the transformer layers. We apply SpecAugment [51] as data augmentation, employing 2 frequency masks (maximum width 30) and 10 time masks (maximum width 50, ratio 0.1). Training batches dynamically combine audio segments up to a total of 7 hours, with gradient clipping set at 0.1. The learning rate (LR) schedule comprises a warm-up phase of 64k steps reaching a peak LR of 0.001, followed by cosine decay for the next 60k steps, and a step decay at a factor of 0.5 thereafter. For MoE ASR models, we include an auxiliary load balancing loss $\mathcal{L}_{\mathrm{load}}$ to encourage even expert utilization, assigning it a weight of 10 based on its relative scale compared to the CTC loss. All models are optimized using AdamW with a weight decay of 0.01 and trained for up to 1M steps (7 epochs) on 16 H100 GPUs, terminating once the greedy validation word error rate (WER) plateaus. Hyperparameters were first optimized using the baseline dense ASR model (84M parameters) and subsequently held constant for all dense and MoE variants. This approach was motivated by the higher computational cost of tuning MoE models. By tuning the dense model and scaling to MoE, we aimed to validate the generalization capabilities of MoE architectures while enabling resource-efficient optimization. However, for some Switch Transformer models, we reduced the peak LR and adjusted the load balancing loss weight to address observed training instability.
V-D Evaluation
We evaluate the final model performance by computing word error rates (WER) using greedy decoding. Before calculating WER, we apply the Whisper normalizer described in [2] to both ground truth and predicted transcriptions.
TABLE I: WERs of a dense ASR, a 2-expert Switch Transformer ASR and a 2-expert Omni-router Transformer ASR for different OOD test datasets.

### v3

# envname-P envname#1

@@ Abstract
[…]
1 Introduction
Automatic speech recognition (ASR) systems have made remarkable progress in recent years, driven by advances in neural network architectures and the availability of large-scale datasets (Vaswani et al.,, 2017; Radford et al.,, 2023; Gulati et al.,, 2020; Kahn et al.,, 2020). Despite these advancements, achieving high recognition accuracy across diverse acoustic and linguistic conditions remains a challenging problem, particularly given the variability in speaker accents, background noise, and domain-specific vocabularies.
There are several ways to improve ASR performance. Combining an acoustic model with additional language models (LMs), such as $n$-gram LMs (Amodei et al.,, 2016; Synnaeve et al.,, 2020), neural LMs (Amodei et al.,, 2016; Synnaeve et al.,, 2020; Zeyer et al.,, 2018; Zeghidour et al.,, 2018) and error correction LMs (Guo et al.,, 2019; Gu et al.,, 2024), were found to be effective ways to produce better transcriptions. Other approaches are inspired by the development of the scaling laws (Brown et al.,, 2020; Hoffmann et al.,, 2022; Chen et al.,, 2025), where scaling up model and data size is crucial to improve model performance. However, classical large models often face trade-offs between computational efficiency and representational power, especially when deployed in real-world scenarios where resource constraints and robustness are critical considerations (He et al.,, 2024; Dai et al.,, 2024).
Mixture-of-experts (MoE) models (Shazeer et al.,, 2017) have emerged as a promising architecture for their ability to scale model capacity (in terms of number of parameters) without incurring proportional overhead at inference. Unlike traditional dense models, which activate all parameters for every input, MoE models dynamically route input data to a subset of specialized experts tailored to process particular characteristics of the data. This routing mechanism enables MoE models to scale to significantly larger parameter counts without proportional increases in inference cost, offering both efficiency and flexibility. ASR involves highly diverse speech signals, with differences in pronunciation, accent, and noise patterns and MoE’s, by design, can be quite appropriate to modeling such diversity. By assigning specific experts to different aspects of the input, MoE models could better capture this variability, potentially improving recognition performance and robustness. Moreover, their ability to adaptively allocate resources makes them well-suited for deployment in heterogeneous environments, from edge devices with constrained resources to cloud-based systems with ample inference capacity. Recently, (Jelassi et al.,, 2024) showed that increasing the number of experts (while fixing the number of active parameters) increases the memorization performance consistently while the reasoning capabilities saturate; (Abnar et al.,, 2025) showed interaction between compute per example and total number of parameters, and their impact on model in MoEs: there is an optimal level of sparsity that improves both training efficiency and model performance.
@@ 1 Introduction
The effectiveness of MoE architectures heavily relies on the routing mechanism that directs inputs to appropriate experts. Conventional MoE methods learn both the routers and expert parameters at the same time, with stochastic gradient descent (SGD). The learning is complicated by issues such as uneven expert utilization (Zhou et al.,, 2022), insufficient expert specialization (Dai et al.,, 2024), and high routing computational costs (Dai et al.,, 2022) often resulting in training instability (Zoph et al.,, 2022). While these issues are well explored, in this paper we highlight an under-explored issue - namely that the correlation between which experts are used at different layers is weak, with different layers making seemly arbitrary, independent decisions about which experts to use. We hypothesize that this would lead to models that do not specialize to data very strongly. To address this issue we propose to use a shared router across all layers that encourages coordinated decision-making across MoE layers (see architecture in Fig. 1). We think it is possible to have a shared router across different layers, because of the residual nature of transformers - features from each layer are modified by residual connections and only change minimally from layer to layer (Sun et al.,, 2025), and thus sharing a decision boundaries across layers is possible by just using a shared router. Empirical results across multiple model configurations consistently indicate that Omni-router delivers superior lower training error, test word error rates (WER) and robustness against highly diverse data compared to traditional dense models and Switch Transformer baselines across diverse out-of-domain evaluation datasets.
2 Related work
@@
MoE models were first introduced by Jacobs et al. (Jacobs et al.,, 1991) as a dynamic mechanism to combine multiple specialized sub-models through gating functions, with subsequent hierarchical extensions presented by Jordan and Jacobs (Jordan and Jacobs,, 1994). Modern advances have significantly scaled MoE architectures to efficiently accommodate extremely large model sizes, exemplified by sparsely gated MoE layers proposed by Shazeer et al (Shazeer et al.,, 2017). Further efficiency improvements were realized with the Switch Transformer (Fedus et al.,, 2022), simplifying the routing mechanism by activating only one expert per input token. Additional strategies, such as BASE layers (Lewis et al.,, 2021) and hash-based routing techniques (Roller et al.,, 2021), have been explored to mitigate issues related to load imbalance and training inefficiency. MoE architectures have demonstrated substantial success across various domains, including natural language modeling (e.g., GLAM (Du et al.,, 2022)) and computer vision tasks (Riquelme et al.,, 2021).
@@
In contrast, MoE applications within ASR remain comparatively limited. SpeechMoE (You et al.,, 2021) proposes a shared embedding network and concatenates the shared embedding with output of the previous layer as the input of routers, meanwhile adding more loss terms, such as $L_{1}$, mean importance and embedding losses. Building on this work, SpeechMoE2 (You et al.,, 2022) proposes a new router architecture by integrating additional global domain and accent embedding into router input to promote adaptability. Both SpeechMoE and SpeechMoE2 predominantly focus on monolingual speech recognition tasks.
@@
Recently, the application of MoE architectures to multilingual ASR has garnered increasing attention (Hu et al.,, 2023; Kwon and Chung,, 2023; Li et al.,, 2024). Hu et al (Hu et al.,, 2023) integrate MoE layers within conformer architectures (Gulati et al.,, 2020), demonstrating an average relative improvement in WER of 11.9% across twelve languages. Kwon et al. (Kwon and Chung,, 2023) propose a MoE model consisting of language-specific experts alongside a language-agnostic expert, while Li et al. (Li et al.,, 2024) utilize routers that select optimal expert subsets combined with joint Connectionist Temporal Classification (CTC)/Attention training (Kim et al.,, 2017) and rescoring, achieving substantial improvements in low-resource multilingual ASR scenarios.
@@
Despite these advancements, critical challenges persist, particularly concerning load balancing, model stability, and training robustness (Zoph et al.,, 2022). Such issues motivate the exploration of novel routing mechanisms aimed at enhancing MoE model robustness and effectiveness. Our contributions address these gaps in two primary ways. First, we introduce an innovative shared-routing approach within Transformer-based MoE ASR models, demonstrating its effectiveness in terms of both accuracy and robustness compared to conventional dense and Switch Transformer MoE architectures. Second, our proposed MoE design remains straightforward, eliminating the need for auxiliary embedding networks, complex auxiliary loss combinations (You et al.,, 2021, 2022), or explicit inductive biases such as accent or language embeddings (You et al.,, 2022; Kwon and Chung,, 2023), thus facilitating simpler implementation and training while being effective.
@@
3 Background
3.1 Mixture-of-Experts layer
Most works consider MoE in replacement of the feed-forward network (FFN) layer in transformer models (Vaswani et al.,, 2017). Typically, an MoE layer contains a routing network and a branch of experts, which are also FFNs. The router dispatches the incoming $T$ tokens $X^{l}\in\mathbb{R}^{T\times D}$ at layer $l$ to different experts first by estimating the weights $P\in\mathbb{R}^{T\times N}$ for each of $N$ experts:
$P^{l}(X^{l})=\mathrm{Softmax}(X^{l}W^{l}),\qquad l=1,...,L,$ (3.1)
@@ 3.1 Mixture-of-Experts layer
3.2 Switch Transformer
Instead of choosing top-$k$ experts based on the router’s decision, Switch Transformer (Fedus et al.,, 2022) simplifies the routing mechanism and chooses only the top-1 expert. It was shown that switch layer not only reduces the router computation cost but also gives better performance in language models. Throughout this paper, we also use only the top-1 expert.
3.3 Load balancing loss
During training of MoE models, an auxiliary loss is often introduced to maintain the balance of load (the number of tokens that go through one expert) of each expert. As introduced in (Fedus et al.,, 2022), given $N$ experts and $T$ tokens $X^{l}$ for layer $l$, the fraction of tokens dispatched to each expert $j$, $f^{l}$, and the fraction of router probability allocated for each expert $j$, $\rho^{l}$, are defined as
$f_{j}^{l}=\frac{1}{T}\sum_{i=1}^{T}\mathbbm{1}\{\arg\max_{j^{\prime}}[P^{l}(X^{l})]_{i}^{j^{\prime}}=j\},$ (3.3)
$\rho^{l}_{j}=\frac{1}{T}\sum_{i:\arg\max_{j^{\prime}}[P^{l}(X^{l})]_{i}^{j^{\prime}}=j}[P^{l}(X^{l})]^{j}_{i},$ (3.4)
@@ 3.3 Load balancing loss
$\mathcal{L}^{l}_{\mathrm{load}}=N\sum_{j=1}^{N}f_{j}^{l}\cdot\rho_{j}^{l}.$ (3.5)
4 Omni-router: sharing routing decisions across MoE layers
@@
$P^{l}(X^{l})=\mathrm{Softmax}(X^{l}W^{\mathrm{shared}}).$ (4.1)
This strategy is viable due to the prevalent use of pre-layer normalization in transformer architectures, which incorporates residual connections directly linking input and output representations. Consequently, feature representations between consecutive layers remain sufficiently similar, enabling shared routing parameters to establish coherent—though not identical—decision boundaries across layers. Meanwhile, the dense parameters within each MoE layer remain free to specialize independently, adapting to layer-specific decision boundaries. This approach contrasts with existing parameter-sharing methods in MoE architectures, which typically involve sharing the experts’ weights rather than the routing parameters themselves (Csordás et al.,, 2024; Zhang et al.,, 2024; Zhao et al.,, 2024).
5 Empirical setup
5.1 Data
We utilize a large-scale conversational audio dataset collected from publicly accessible sources, named SpeechCrawl. This dataset was specifically curated to be diverse, conversational, multilingual, and multi-speaker. Audio samples average approximately 30 minutes in duration, with around 60% consisting of English speech. Since SpeechCrawl does not provide transcriptions, we generate pseudo-labels using the WhisperX (Bain et al.,, 2023) pipeline in combination with Whisper’s large-v2 multilingual model (Radford et al.,, 2023). These pseudo-labels simultaneously yield audio segmentation (with segments shorter than 30 seconds) and corresponding transcripts. To ensure quality, we filter out unreliable pseudo-labels resulting from looping artifacts of decoding in the sequence-to-sequence model. Specifically, we employ a simple heuristic based on expected speech rates (approximately three words per second), similar to (Likhomanenko et al.,, 2021). After filtering, we retain approximately 1M hours of English speech segments for training. Additionally, we randomly select about 12h of audio (24 files) as our development set to monitor training progress and identify the optimal checkpoint for subsequent evaluation on out-of-domain (OOD) datasets. For OOD testing, we employ diverse test sets from AMI-IHM (Carletta,, 2007), Callhome (Godfrey et al.,, 1992), Chime6 (Watanabe et al.,, 2020), CommonVoice (Ardila et al.,, 2019), Fleurs (Conneau et al.,, 2023), Librispeech (Panayotov et al.,, 2015), Switchboard (Godfrey et al.,, 1992), Tedlium (Hernandez et al.,, 2018), Voxpopuli (Wang et al.,, 2021), and Wall Street Journal (Paul and Baker,, 1992). Importantly, none of the selected test datasets overlap with the training data.
5.2 ASR models
We compare our proposed Omni-router Transformer model against two baseline architectures: a dense transformer and a Switch Transformer. All ASR models utilize the standard Transformer encoder architecture (Vaswani et al.,, 2017), optimized using CTC (Graves and Graves,, 2012) and employing an 8k-token word-piece vocabulary derived from the training transcriptions.
@@ 5.2 ASR models
5.3 Training setup
All ASR models take as input 80-dimensional log-mel filterbank features computed using a 25ms sliding window with a 10ms stride. Subsequently, every four consecutive frames are stacked before being fed into the transformer layers. We apply SpecAugment (Park et al.,, 2019) as data augmentation, employing 2 frequency masks (maximum width 30) and 10 time masks (maximum width 50, ratio 0.1). Training batches dynamically combine audio segments up to a total of 7 hours, with gradient clipping set at 0.1. The learning rate (LR) schedule comprises a warm-up phase of 64k steps reaching a peak LR of 0.001, followed by cosine decay for the next 60k steps, and a step decay at a factor of 0.5 thereafter. For MoE ASR models, we include an auxiliary load balancing loss $\mathcal{L}_{\mathrm{load}}$ to encourage even expert utilization, assigning it a weight of 10 based on its relative scale compared to the CTC loss. All models are optimized using AdamW with a weight decay of 0.01 and trained for up to 1M steps (7 epochs) on 16 H100 GPUs, terminating once the greedy validation word error rate (WER) plateaus. Hyperparameters were first optimized using the baseline dense ASR model (84M parameters) and subsequently held constant for all dense and MoE variants. This approach was motivated by the higher computational cost of tuning MoE models. By tuning the dense model and scaling to MoE, we aimed to validate the generalization capabilities of MoE architectures while enabling resource-efficient optimization. However, for some Switch Transformer models, we reduced the peak LR and adjusted the load balancing loss weight to address observed training instability.
5.4 Evaluation
We evaluate the final model performance by computing word error rates (WER) using greedy decoding. Before calculating WER, we apply the Whisper normalizer described in (Radford et al.,, 2023) to both ground truth and predicted transcriptions.
6 Results
6.1 Omni-router improves ASR
In Table 1, we present the WER results comparing a baseline dense ASR model (140M) with Switch Transformer and Omni-router Transformer ASR models, each employing two experts per layer. We first note that both MoE-based architectures consistently outperform the dense baseline across all evaluated out-of-domain datasets. Additionally, the Omni-router Transformer further improves upon the performance of the Switch Transformer. Given that the Omni-router shares hyperparameters originally optimized for the Switch Transformer, its superior performance despite minimal hyperparameter tuning underscores the effectiveness and robustness of the shared-routing strategy.
Table 1: WERs of a dense ASR, a 2-expert Switch Transformer ASR and a 2-expert Omni-router Transformer ASR for different OOD test datasets.